Integrate the function with respect to X
1/(x-1)(x+2) by partial fraction

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
1/(x-1)(x+2) = A/(x-1) +( Bx + C)/(x-1)^2 + D/(x + 2)
 
Now 1 = A(x^2 + x – 2) + Bx (x+2) + C(x+2) + D (x^2 + 1 – 2x)
 
Now compare and find A, B ,C  And D.